Problem 7-13 Owen Conner works part-time packaging software for a local distribution company in Indiana. The annual fixed cost is $15,000 for this process, direct labor is $4.00 per package, and material is $4.75 per package. The selling price will be $12.50 per package 10 points a. What is the break-even point in units? (Do not round intermediate calculations. Roundup your answer to the next whole number) units Break-even point eBook Hint References b. How much revenue do we need to take in before breaking even? (Round your answer to the nearest dollar amount.) ak-even point
Answer #1

Breakeven point is defined as the point where the total revenue is equal to the total cost. Here total cost means the fixed cost and the variable cost involved. If number of packages produced is x to reach the breakeven point, then the following equation holds true at breakeven point

S*x = FC + x* (DL+M)

where S is the selling price per package, FC is the fixed cost, DL is the direct labour cost per package and M is the material cost per package

For the given problem

12.5*x = 15000 + x*(4 + 4.75)

12.5*x = 15000 + 8.75*x

12.5*x - 8.75*x = 15000

3.75*x = 15000

x = \frac{15000}{3.75}

x = 4000

Hence the breakeven points in units is 4000

The revenue to be taken in before breaking even is given by breakeven units multiplied by the selling price. Hence breakeven revenue is equal to 12.5*4000 = $ 50000